OpenStudy (anonymous):

a unit vector in the opposite direction to 4i-j-k

OpenStudy (mathmate):

A vector's opposite has the sign of all the components reversed.

OpenStudy (nirmalnema):

-4i+j+k

OpenStudy (anonymous):

so the answer will be \[\frac{ 1 }{ 3\sqrt{2} }\] (-4+j+k) or \[\frac{ 1 }{ \sqrt{14} }\](-4+j+k)

OpenStudy (anonymous):

hello people unit vector has got magnitude 1 .. is ur answwer giving magnitude 1 ??

OpenStudy (anonymous):

yes need magnitude.

hartnn (hartnn):

divide by magnitude to get the unit vector magnitude of \(xi+yj+zk\) will be \(\sqrt{x^2+y^2+z^2}\)

hartnn (hartnn):

\(\Large \sqrt{(-4)^2+1^2+1^2} = ...\)

OpenStudy (anonymous):

thank u!! i got the answer :)

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

since you seem to be new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

thank you soo much!! Its so kind of you :)